Given the following system of equations, identify the type of system.

y = -x + 4
2x + 2y = 8

Answer by ikleyn(52784)

.
Given the following system of equations, identify the type of system.
     y = -x + 4
    2x + 2y = 8
----------------

The starting equations are

     y = -x + 4
    2x + 2y = 8


Write equations in standard form

    x +  y = 4
   2x + 3y = 8


You see that the second equation is the doubled the first one.


In other words, two equations of the system are equivalent.


It is the same as if the system consists of one equation in two unknowns.


Such a system of equations is consistent and has infinitely many solutions.

Solved, answered and explained.

------------------

See the lesson
    - Geometric interpretation of the linear system of two equations in two unknowns
in this site.